OpenStudy (anonymous):

plz help. I need someone to check my work.. is this right 5x^3/4y^4?

jimthompson5910 (jim_thompson5910):

I'm guessing 5x^3/4y^4 is your answer?

jimthompson5910 (jim_thompson5910):

If so, what's the question?

OpenStudy (anonymous):

srry I have to pu load the pic.

jimthompson5910 (jim_thompson5910):

ok

jimthompson5910 (jim_thompson5910):

one second

OpenStudy (anonymous):

okay. :)

jimthompson5910 (jim_thompson5910):

unfortunately what you got is incorrect

OpenStudy (anonymous):

is it this? I'm lost..5x^4/4y^3??

jimthompson5910 (jim_thompson5910):

no it is not

OpenStudy (anonymous):

can you show me how to do it?

jimthompson5910 (jim_thompson5910):

\[\Large \frac{15x^5}{24y^8} \div \frac{8y^4}{4x^2}\] \[\Large \frac{15x^5}{24y^8} \times \frac{4x^2}{8y^4}\] Did you get this far?

OpenStudy (anonymous):

okay gottcha.

jimthompson5910 (jim_thompson5910):

then what's next?

OpenStudy (anonymous):

you cross multipy. but 24 can go into 8 ...3 times. but then I get lost..

jimthompson5910 (jim_thompson5910):

Well you multiply straight across, then you reduce as much as possible \[\Large \frac{15x^5}{24y^8} \div \frac{8y^4}{4x^2}\] \[\Large \frac{15x^5}{24y^8} \times \frac{4x^2}{8y^4}\] \[\Large \frac{15x^5 \times 4x^2}{24y^8 \times 8y^4}\] \[\Large \frac{60x^7}{192y^{12}}\] \[\Large \frac{12*5x^7}{12*16y^{12}}\] \[\Large \frac{\cancel{12}*5x^7}{\cancel{12}*16y^{12}}\] \[\Large \frac{5x^7}{16y^{12}}\]

jimthompson5910 (jim_thompson5910):

So the final answer is \[\Large \frac{5x^7}{16y^{12}}\]

OpenStudy (anonymous):

I get how your breaking it down , But that's not one of my choices?

jimthompson5910 (jim_thompson5910):

what are your choices?

OpenStudy (anonymous):

1min

jimthompson5910 (jim_thompson5910):

ok

jimthompson5910 (jim_thompson5910):

is the problem really \[\Large \frac{15x^5}{24y^8} \times\frac{8y^4}{4x^2}\]

jimthompson5910 (jim_thompson5910):

and NOT with a division sign in the middle?

OpenStudy (anonymous):

it with the divison sign

OpenStudy (anonymous):

it's*

jimthompson5910 (jim_thompson5910):

weird

jimthompson5910 (jim_thompson5910):

i see the answer if it had a multiplication symbol and not a division symbol

OpenStudy (anonymous):

you have it back words the last part. it shud be 4x^2/8y^4

jimthompson5910 (jim_thompson5910):

So the actual problem is \[\Large \frac{15x^5}{24y^8} \div\frac{4x^2}{8y^4}\] ???

OpenStudy (anonymous):

yes.

jimthompson5910 (jim_thompson5910):

ok that makes a ton more sense

jimthompson5910 (jim_thompson5910):

\[\Large \frac{15x^5}{24y^8} \div \frac{4x^2}{8y^4}\] \[\Large \frac{15x^5}{24y^8} \times \frac{8y^4}{4x^2}\] \[\Large \frac{15x^5 \times 8y^4}{24y^8 \times 4x^2}\] \[\Large \frac{120x^5y^4}{96x^2y^8}\] \[\Large \frac{24*5x^5y^4}{24*4x^2y^8}\] \[\Large \frac{\cancel{24}*5x^5y^4}{\cancel{24}*4x^2y^8}\] \[\Large \frac{5x^5y^4}{4x^2y^8}\] \[\Large \frac{5x^{5-2}y^{4-8}}{4}\] \[\Large \frac{5x^3y^{-4}}{4}\] \[\Large \frac{5x^3}{4y^{4}}\]

OpenStudy (anonymous):

thank you very much, your a big help

jimthompson5910 (jim_thompson5910):

i'm glad, yw